How do you solve log4(9z+5)-log4(2z+2)=1? I've went over all of my lessons but I still don't understand it. Please show all work so I understand how to do it.
Answer by lwsshak3(11628)

How do you solve log4(9z+5)-log4(2z+2)=1? I've went over all of my lessons but I still don't understand it. Please show all work so I understand how to do it.
**
log4(9z+5)-log4(2z+2)=1
use division rule for logs
log4[(9z+5)/2z+2)]=1
convert to exponential form: base(4) raised to log of the number(1)=number(9z+5)/2z+2)
4^1=(9z+5)/2z+2)
4=(9z+5)/2z+2)
9z+5=8z+8
z=3